OpenStudy (trojanpoem):

w, w^2 , 1 are the roots of the 1 w^w = .....

OpenStudy (phi):

if w, w^2 , 1 are all of the roots, then w is the cubed root of 1 if we write 1 in exponential form \[ e^{2 \pi i} = 1\] then w is \[ w = e^{\frac{2 \pi}{3} i }\] in rectangular a+bi form this is \[ w = -\frac{1}{2} + \frac{\sqrt{3}}{2} i \] thus we can write \[ w^w= \left( e^{\frac{2 \pi}{3} i }\right)^{ -\frac{1}{2} + \frac{\sqrt{3}}{2} i } \] use \[ \left(a^b\right)^c = a^{bc} \] so simplify

OpenStudy (trojanpoem):

w^w = e^(-pi/3 * i - pi/sqrt(3)) How will I get rid of this i ? the question is asking for a value.

OpenStudy (phi):

the answer is complex you could write it in the form \[ A \exp ( i \theta) \] or \[ A\cos(\theta) + i A\sin(\theta) \] here A is \( \exp( -\pi/\sqrt{3}) \) and theta = -pi/3

OpenStudy (trojanpoem):

Thanks a lot.